# Audio Flamingo: A Novel Audio Language Model with Few-Shot Learning and Dialogue Abilities

### v2

@@ B.3 Dialogue filtering
The dialogues generated by GPT-4 as discussed in Section B.2 do not always follow the prompts, resulting in answers that have phrases such as “does not specify”, “cannot be determined”, “without additional context” and so on. Hence, following Gardner et al. (2023), we filter such QA pairs to improve the data quality and ensure desirable outputs from the model. Apart from the manual filtering step, we also filter samples based on the similarity of the answer generated by GPT-4 and the audio samples. Specifically, we compute the cosine similarity between the LAION-CLAP text-embeddings and audio-embeddings (Wu et al., 2023) for a given QA pair in each dialogue. The distributions of similarities are shown in Figure 6. We remove samples if the similarity is below 0.3.
Figure 6: Distributions of LAION-CLAP similarities before filtering.
@@ B.4 Examples of generated dialogues
[Annotations: Sound of Mechanisms (Produces sound through mechanical means.): [0.119s-10.000s]; Sound of Generic impact sounds (Short, sharp, and loud bursts.): [0.217s-0.387s], [2.394s-2.727s], [2.889s-3.255s], [3.799s-4.132s], [4.473s-4.766s], [4.928s-5.188s], [8.129s-10.000s]; Sound of Glass shatter (High-pitched, sharp, sudden, loud crash.): [1.622s-4.652s]; Sound of Walk, footsteps (Soft, rhythmic, echoing thuds.): [5.724s-6.106s], [6.455s-6.650s], [7.024s-7.479s], [7.625s-7.901s]] User: Can you describe the main events in the audio? Assistant: The audio mainly consists of mechanical sounds, sharp impact sounds, the shattering of glass, and the sound of footsteps. User: What can you tell me about the impact sounds? Assistant: The impact sounds are short, sharp, and loud bursts. They occur at several points throughout the audio, suggesting some kind of repetitive action or event. User: What about the glass shattering, does it happen more than once? Assistant: The glass shattering sound is a high-pitched, sharp, sudden, loud crash. It occurs once in the audio, lasting for about three seconds.
[Annotations: This is the recording of a cowbell. The cowbell is being played with the side and the tip of the drumstick repeatedly. This recording could be sampled for use in beat-making especially for genres that have a heavy usage of cowbells such as latin dance music.] User: What instrument is being played in this audio? Assistant: The instrument being played in this audio is a cowbell. User: How is it being played? Assistant: The cowbell is being played with the side and the tip of the drumstick repeatedly. User: What is the tone of the cowbell? Assistant: The tone of the cowbell is bright, resonant, and penetrating.
Appendix C Additional Examples and Discussion of Audio Flamingo
@@ C.1 Additional Few-shot Experiments on Unseen Labels
The labels for BG-Gun-Sound are:
No gun, SLR, M4, VSS, AK, SKS, PP, P90, AUG, SCAR, UMP, QBZ, P92, M249, QBU, DEAGLE, TOMY, PUMP, P1911, K2, DP, GRO, AWM, VEC, UZI, M24, KAR, VERL, MINI, WIN, MK, R45, M16, P118C, G36C, S12K, DBS, R1895.
The labels for Birds-Dataset-Till-2022 are:
Brazilian Tinamou, Red-legged Tinamou, Hooded Tinamou, Chestnut-bellied Guan, Dwarf Cassowary, Little Spotted Kiwi, Dwarf Tinamou, Baudo Guan, Little Chachalaca, Pale-browed Tinamou, Lesser Rhea, Australian Brushturkey, Berlepsch’s Tinamou, Spotted Nothura, Puna Tinamou, Brushland Tinamou, Andean Guan, Black-billed Brushturkey, Blue-throated Piping Guan, Somali Ostrich, Wattled Brushturkey, White-bellied Chachalaca, Barred Tinamou, Chaco Chachalaca, Rufous-headed Chachalaca, Plain Chachalaca, Emu, Southern Cassowary, White-bellied Nothura, Undulated Tinamou, Tawny-breasted Tinamou, Thicket Tinamou, Bearded Guan, Quebracho Crested Tinamou, Scaled Chachalaca, Micronesian Megapode, Nicobar Megapode, White-throated Tinamou, Malleefowl, Crested Guan, Chestnut-winged Chachalaca, Speckled Chachalaca, Ornate Tinamou, Curve-billed Tinamou, Tataupa Tinamou, White-crested Guan, Southern Brown Kiwi, Sula Megapode, Biak Scrubfowl, Vanuatu Megapode, Patagonian Tinamou, Moluccan Megapode, Variegated Tinamou, White-browed Guan, Marail Guan, Chestnut-headed Chachalaca, Northern Cassowary, Bartlett’s Tinamou, Rusty-margined Guan, Spix’s Guan, Band-tailed Guan, Andean Tinamou, Trinidad Piping Guan, Elegant Crested Tinamou, North Island Brown Kiwi, Cinereous Tinamou, White-winged Guan, Grey-headed Chachalaca, Okarito Kiwi, Rufous-bellied Chachalaca, Red-winged Tinamou, Rufous-vented Chachalaca, Black-fronted Piping Guan, Tanimbar Megapode, Choco Tinamou, Great Spotted Kiwi, Little Tinamou, Slaty-breasted Tinamou, Tepui Tinamou, Greater Rhea, Chilean Tinamou, Cauca Guan, Yellow-legged Tinamou, Black Tinamou, Lesser Nothura, Rusty Tinamou, Grey-legged Tinamou, Dusky Megapode, Red-billed Brushturkey, Dusky-legged Guan, Collared Brushturkey, Common Ostrich, Colombian Chachalaca, Melanesian Megapode, Maleo, New Guinea Scrubfowl, Red-throated Piping Guan, Huayco Tinamou, Buff-browed Chachalaca, Black-capped Tinamou, Philippine Megapode, Great Tinamou, Orange-footed Scrubfowl, Brown Tinamou, Solitary Tinamou, Taczanowski’s Tinamou, Grey Tinamou, Small-billed Tinamou, Tongan Megapode, Highland Tinamou, West Mexican Chachalaca, East Brazilian Chachalaca, Red-faced Guan, Darwin’s Nothura.
C.2 Audio Flamingo understands sound in depth
@@
C.3 Audio Flamingo looks at retrieved samples
@@
C.4 Audio Flamingo has strong dialogue ability
@@
[Audio annotations: Sound of Music (Sound produced by vibrating instruments.): [0.000s-10.000s]; Sound of Male singing (Deep, resonant, and powerful tones.): [1.341s-3.734s], [9.289s-10.000s]] User: What genre does this music belong to? Ground Truth: This music belongs to the R&B/Disco genre. Baseline: It is not specified in the given audio clip. Audio Flamingo: This music belongs to the R&B genre. User: Can you tell me more about the vocals in this track? Ground Truth: Sure, the vocals in this track are from a male singer who is singing in a higher register. His voice has been modified with effects, making it sound like a voice sample. Baseline: The vocals in this track are described as rich and full-bodied, but no further details are given about the specific singer or the lyrics being sung in the audio clip. Audio Flamingo: The vocals in this track are by a male singer. His voice is deep, resonant, and powerful, adding to the overall emotional impact of the song.

### v3

@@ B.3 Dialogue filtering
The dialogues generated by GPT-4 as discussed in B.2 do not always follow the prompts, resulting in answers that have phrases such as “does not specify”, “cannot be determined”, “without additional context” and so on. Hence, following Gardner et al. (2023), we filter such QA pairs to improve the data quality and ensure desirable outputs from the model.
Apart from the manual filtering step, we also filter samples based on the similarity of the answer generated by GPT-4 and the audio samples. Specifically, we compute the cosine similarity between the LAION-CLAP text-embeddings and audio-embeddings (Wu et al., 2023) for a given QA pair in each dialogue. The distributions of similarities are shown in Figure 6. We remove samples if the similarity is below a threshold of 0.3.
Figure 6: Distributions of LAION-CLAP similarities before filtering the generated AF-Dialogue-AudioSetSL (left) and AF-Dialogue-MusicCaps (right) datasets.
@@ B.4 Examples of generated dialogues
We demonstrate an example of generated dialogue for each of the generated dataset.
[Audio annotations: Sound of Mechanisms (Produces sound through mechanical means.): [0.119s-10.000s]; Sound of Generic impact sounds (Short, sharp, and loud bursts.): [0.217s-0.387s], [2.394s-2.727s], [2.889s-3.255s], [3.799s-4.132s], [4.473s-4.766s], [4.928s-5.188s], [8.129s-10.000s]; Sound of Glass shatter (High-pitched, sharp, sudden, loud crash.): [1.622s-4.652s]; Sound of Walk, footsteps (Soft, rhythmic, echoing thuds.): [5.724s-6.106s], [6.455s-6.650s], [7.024s-7.479s], [7.625s-7.901s]] User: Can you describe the main events in the audio? Assistant: The audio mainly consists of mechanical sounds, sharp impact sounds, the shattering of glass, and the sound of footsteps. User: What can you tell me about the impact sounds? Assistant: The impact sounds are short, sharp, and loud bursts. They occur at several points throughout the audio, suggesting some kind of repetitive action or event. User: What about the glass shattering, does it happen more than once? Assistant: The glass shattering sound is a high-pitched, sharp, sudden, loud crash. It occurs once in the audio, lasting for about three seconds.
[Audio annotations: This is the recording of a cowbell. The cowbell is being played with the side and the tip of the drumstick repeatedly. This recording could be sampled for use in beat-making especially for genres that have a heavy usage of cowbells such as latin dance music.] User: What instrument is being played in this audio? Assistant: The instrument being played in this audio is a cowbell. User: How is it being played? Assistant: The cowbell is being played with the side and the tip of the drumstick repeatedly. User: What is the tone of the cowbell? Assistant: The tone of the cowbell is bright, resonant, and penetrating.
Appendix C Additional Experiments, Examples and Discussion of Audio Flamingo
@@ C.1 Additional Few-shot Experiments on Unseen Labels
C.2 Compute
In this section we compare the computation resources required by our model and baseline models. Pengi (Deshmukh et al., 2023) is based on GPT2-base and has a total size 308M. It’s trained on 20 V100 GPUs. LTU (Gong et al., 2023c) is based on LLaMA-7B and has a total size 6.74B. It’s trained on 4 RTX A6000 GPUs for about 3 days. Qwen-Audio (Chu et al., 2023) is based on Qwen-7B and has a total size of 8.3B. The training cost is unknown. Salmonn (Tang et al., 2023a) is based on Vicuna-13B. The training cost is unknown. Our proposed Audio Flamingo has a total size of 2.2B. We use 8 A100 GPUs. It takes about 5-6 days for pretraining, about 3 days for SFT, and about 2hrs for dialogue fine-tuning.
C.3 Audio Flamingo understands sound in depth
@@
C.4 Audio Flamingo looks at retrieved samples
@@
C.5 Audio Flamingo has strong dialogue ability
@@
[Audio annotations: Sound of Music (Sound produced by vibrating instruments.): [0.000s-10.000s]; Sound of Male singing (Deep, resonant, and powerful tones.): [1.341s-3.734s], [9.289s-10.000s], ] User: What genre does this music belong to? Ground Truth: This music belongs to the R&B/Disco genre. Baseline: It is not specified in the given audio clip. Audio Flamingo: This music belongs to the R&B genre. User: Can you tell me more about the vocals in this track? Ground Truth: Sure, the vocals in this track are from a male singer who is singing in a higher register. His voice has been modified with effects, making it sound like a voice sample. Baseline: The vocals in this track are described as rich and full-bodied, but no further details are given about the specific singer or the lyrics being sung in the audio clip. Audio Flamingo: The vocals in this track are by a male singer. His voice is deep, resonant, and powerful, adding to the overall emotional impact of the song.